a bullet from a 0.38 caliber handgun having a 6 inch barrel attains a muzzle speed of meters per second. assuming constant acceleration, how much time does it take the bullet to travel down the barrel?
1
Expert's answer
2012-03-01T11:04:50-0500
We know that at²/2 = L, where L is a barrel length, a is an acceleration and t is a time while bullet is in a barrel.
Also we know that at = V, here V is a muzle speed. So, a = V/t and
at²/2 = L rewrites as
(V/t)*t²/2 = L, or
Vt/2 = L ==> t = 2L/V = 2*6/V = 12/V.